OpenStudy (anonymous):

Solve 11c – 2(2c + 1) = 3(c – 6)

OpenStudy (unklerhaukus):

\[11c – 2(2c + 1) = 3(c – 6)\]expand the brackets\[11c-4c-2=3c-18\]simplify the c's on the left hand side\[7c-2=3c-18\]-3c and add to to both sides\[4c=16\]divide by 4\[c=4\]

OpenStudy (unklerhaukus):

*-3c and add 2 to both sides